2{[3(y-4)+19]-[2(5y-3)+2]}
Answer by tommyt3rd(5050)

2{[3(y-4)+19]-[2(5y-3)+2]}=
2{[3y-12+19]-[10y-6+2]}=
2{3y+7-[10y-4]}=
2{3y+7-10y+4}=
2(-7y+11)=
-14y+22